The difference between the years of birth and death of a lady is 40 year two times the year of her birth exceed the year of her deathy by 335. find the year of her birth
Answer by ankor@dixie-net.com(22740)

The difference between the years of birth and death of a lady is 40 years.
Let x = her birth year
then
(x+40) = her death year
:
two times the year of her birth exceed the year of her death by 335.
find the year of her birth
2x = (x+40) + 335
2x - x = 40 + 335
x = 375 is the year of her birth